OpenStudy (anonymous):

Find the 7th partial sum of summation of 6 open parentheses 3 close parentheses to the I minus 1 power from 1 to infinity

OpenStudy (anonymous):

\[\sum_{i=1}^{\infty} 6(3)^i-1\]

OpenStudy (anonymous):

it's suppose to be 6(3)^i-1

OpenStudy (anonymous):

2,184 6,558 27,993 167,961

OpenStudy (anonymous):

I got B :)

jimthompson5910 (jim_thompson5910):

Me too Sn = a*(1-r^n)/(1-r) S7 = 6*(1-3^7)/(1-3) S7 = 6,558

OpenStudy (anonymous):

Yayyy! Twinning hahah :D Thanks

jimthompson5910 (jim_thompson5910):

no problem